Assume there are no transfer payments and taxes and Yd=Y, given the following functional values answer the questions below:

Consumption function: C=100+0.7YD

Planned investment: I=40

Government expenditure: G=80

Find the equilibrium level of income in the economy.                        

Find the equilibrium level of income in the economy if government expenditure increases to 120 and find the change in the equilibrium income level from part a to b.                           



1
Expert's answer
2021-06-03T13:44:22-0400

"Z=C+I+G\\Z=100+0.7Y_D+40+80"


"Y_D=Y"

so, "Z=100+0.7Y+40+80"


The equilibrium condition is that aggregate demand is equal to output

"Z=Y"


"Y=100+0.7Y+40+80\\Y=220+0.7Y\\Y-0.7Y=220\\Y=733.33"


When the level of income in economy increases to 120,

"Y=100+0.7Y+40+120\\Y=260+0.7Y\\0.3Y=260\\Y=866.67"


Percentage change

"\frac{866.67-733.33}{733.33}\times 100\%=18.2\%"